OpenStudy (anonymous):

Help Let f(x) = x2 – 8x + 5. Find f(–1). –4 12 13 14 I think its the first one but I am not sure?

jimthompson5910 (jim_thompson5910):

you replace x with -1, then evaluate/simplify

OpenStudy (anonymous):

I got 1?

jimthompson5910 (jim_thompson5910):

not sure how

OpenStudy (anonymous):

Let f(x) = x2 – 8x + 5. Find f(–1). -1(2)-8(-1)+5

OpenStudy (anonymous):

-2-10+5 -2-10=-12+5=-7

jimthompson5910 (jim_thompson5910):

x^2 means x squared

jimthompson5910 (jim_thompson5910):

replace x with -1 to get (-1)^2 = 1

OpenStudy (anonymous):

oh,

jimthompson5910 (jim_thompson5910):

you're not doing x2 = (-1)*2 = -2

jimthompson5910 (jim_thompson5910):

also, -8(-1) = +8

OpenStudy (anonymous):

Let f(x) = x2 – 8x + 5. Find f(–1). -1^+8x+5= 12

OpenStudy (anonymous):

So B?

jimthompson5910 (jim_thompson5910):

f(x) = x^2 – 8x + 5 f(-1) = (-1)^2 – 8(-1) + 5

OpenStudy (anonymous):

1+8+5= 14 So D?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

YAY thank you.

jimthompson5910 (jim_thompson5910):

yw